The probability that a microscope is defective is .17%. Find the probability that one of the first 50 microscopes is defective
1
Expert's answer
2019-10-24T09:12:27-0400

Let "X" be the number of defective microscopes. Because the events (finding that certain microscope is defecrive or not) have equal probability "p = 1.7 \cdot {10^{ - 3}}" that it will be defecrive, we can conclude that "X" follows the binomial distribution "X \sim B(n,p)" (where "n = 50") with probability mass function


"{p_X}(k) = C_n^k{p^k}{(1 - p)^{n - k}}"

where "C_n^k = \frac{{n!}}{{k!(n - k)!}}" is the binomial coefficient.

Under "one of the first" they mean "exactly one", we need to find the probability "P(X = 1)", thus

"P(X = 1) = {p_X}(1) = \frac{{50!}}{{1!49!}} \cdot 1.7 \cdot {10^{ - 3}} \cdot {(1 - 1.7 \cdot {10^{ - 3}})^{49}} \approx 0.0782008"

If they mean "one or more", we need to find the probability "P(X \geqslant 1)" that can be rewritten as


"P(X \geqslant 1) = 1 - P(X < 1) = 1 - P(X = 0)"

and


"P(X = 0) = {p_X}(0) = \frac{{50!}}{{0!50!}} \cdot {(1.7 \cdot {10^{ - 3}})^0} \cdot {(1 - 1.7 \cdot {10^{ - 3}})^{50}} \approx 0.918446"

thus


"P(X \geqslant 1) \approx 1 - 0.918446 \approx 0.081554"

Answer: "P(X = 1) \approx 0.0782008".